# InfoAffect: A Dataset for Affective Analysis of Infographics

### v1

@@ 3.3.1 JSON Format for Affects
The JSON format used in InfoAffect dataset follows a structured key–value schema to represent the affective analysis results of each infographic and its corresponding textual description. Table 3 described the function of each key-value pair in the JSON file. For every infographic–text pair, we provided a single JSON file containing affect-related annotations under the “affective_analysis” key. Each file included a list of affects and their corresponding confidence scores within the “affective_words” array. This design enabled researchers to easily access both the lexical and quantitative dimensions of affective representation.
@@
Table 3: The table provided a detailed explanation of the JSON structure adopted in InfoAffect dataset for organizing affective annotations. Each key–value pair specified how the detected affects, corresponding confidence scores, and relevant lexical information are stored and accessed. This standardized format supported transparent data management, reproducibility of experiments, and seamless integration with downstream analytical workflows.
@@
3.3.2 Prompting Strategies for Affect Extraction
Table 4: This table presented the detailed configuration of the TIDD-EC prompt framework employed for affect extraction. Each component defined a specific role in guiding multimodal large language models, from specifying the analytical task and procedural instructions to providing explicit behavioral constraints and contextual examples.
@@ 3.3.2 Prompting Strategies for Affect Extraction
D-Do Specify the actions that MLLMs should implement during the response process. Provide your analysis in the following JSON structure. Output ONLY a single valid JSON object.
@@
Prompt frameworks have demonstrated their effectiveness in optimizing interactions with MLLMs, thereby achieve more faithful instruction following [45]. In this study, we employed the TIDD-EC prompt framework [47], which was composed of five essential components: task type, detailed instructions, dos and don’ts, examples, and user content. By systematically defining the task scope, providing explicit operational guidelines, specifying behavioral constraints, illustrating desired results through examples, and incorporating user specific contextual information, the TIDD-EC prompt framework contributed to generating responses that are both accurate and relevant. We maintain MLLMs were instructed to select affects exclusively from the Affect Table introduced in Sect. 3.2, rather than producing words freely. Each selected affect was required to be accompanied by a corresponding confidence score. Table 4 summarizes the detailed prompt configurations applied to the affect extraction task under the TIDD-EC prompt framework.
3.3.3 Affective Result Fusion
To fully leverage the complementary strengths of different models, we integrated the results from the five models introduced in the beginning of Sect. 3.3 using the RRF algorithm [7]. RRF algorithm is an effective fusion technique that requires no training and is independent of the initial scores. This fusion strategy has been widely demonstrated to perform robustly across multimodal and retrieval tasks, effectively balancing precision and recall while enhancing the stability and consistency of final affective predictions [50, 41]. The fusion process was explained in Fig. 4.
$\text{s}(A)=\sum_{i=1}^{5}\frac{C_{i(A)}}{k+r_{i(A)}}$ (5)
Table 5: The table details the demographic and occupational composition of participants who took part in the InfoAffect dataset evaluation study. It lists their professional identities, gender, usage scenarios, and age. The group comprises researchers, students, designers, and data professionals with backgrounds in visualization and communication design.
@@ 3.3.3 Affective Result Fusion
4 User Study
In this section, we elaborated on the detailed design, implementation process, participant composition, data collection methods, and result analysis of the practical case study, aiming to verify the usability and accuracy of the InfoAffect dataset. Representative examples and the corresponding analytics for this section were illustrated in Fig. 5.
Figure 5: Visual comparison of CACI results between InfoAffect dataset and crowdsourced ratings. Panels (a–i) were the representative examples cited in this section (labels on each panel); they illustrate typical cases rather than the full dataset. Each row corresponds to one affective category, indicated by colored legend markers: $\blacksquare$ Negative, $\blacksquare$ Neutral, and $\blacksquare$ Positive.
4.1 Participants
We conducted a user study with 12 participants who had expertise in data journalism and visualization. Participants were recruited through Zhejiang University of Technology and local industry partners, which comprised two university faculty members (E1–E2) who teach courses on data visualization, three students (P1–P3), and seven industry practitioners (P4–P10), including data analysts, open-data officers, policy researchers, data consultants, marketing analyst, independent designers and design blogger. To protect privacy in reporting, we refer to participants using these pseudonymous identifiers (E1–E2, P1–P10) throughout the paper.
Table 5 summarizes the participant demographics, including gender, typical usage scenarios, and age. Participants ranged from 26 to 33 years old (mean = 29.3), and the group included five females and seven males, with a balanced gender composition and diverse disciplinary and professional backgrounds. This diversity enabled us to capture representative perspectives across academia and industry, and to compare needs and practices across instructional, student, and professional roles. This user study received approval from the Institutional Review Board (IRB) of our university. Prior to participation, all individuals provided written informed consent. The collected data were used exclusively for research purposes, and strict confidentiality procedures were followed to safeguard participants’ personal information. Each participant received \$15 as compensation for their efforts.
4.2 Exp. I : Evaluating Usability of InfoAffect dataset
4.2.1 Study Procedure
Following the same processing pipeline used in constructing the InfoAffect dataset described in Sec 3.3, we analyzed each infographic to extract its affects. The pipeline automatically extracted affects, which were then compared with the participants’ original intent. After that, We presented the extracted affects to the participants and conducted a semi-structured interview. Questions are as following:
@@ 4.2.1 Study Procedure
  • Do you think the extracted affect aligns with the affect you originally intended to convey in this infographic?
@@
  • If the extracted affects differs from your original intent, how would you describe that difference?
@@
  • Do you think which visual elements contributed to this unintended affects?
@@
  • Does the analysis results help you better understand how affect is represented in your visual design?

### v2

@@ Introduction
The JSON schema used in InfoAffect dataset follows a structured key–value schema to represent the affective analysis results of each infographic and its corresponding textual description. (Appendix C) described the function of each key-value pair in the JSON file. For every infographic–text pair, we provided a single JSON file containing affect-related annotations under the “affective_analysis” key. Each file included a list of affects and their corresponding confidence scores within the “affective_words” array. This design enabled researchers to easily access both the lexical and quantitative dimensions of affective representation.
Prompting Strategies for Affect Extraction
Prompt frameworks have demonstrated their effectiveness in optimizing interactions with MLLMs, thereby achieve more faithful instruction following (Teo 2024; Wang et al. 2025). In this study, we employed the TIDD-EC prompt framework (Vivas.AI 2024), which was composed of five essential components: task type, detailed instructions, does and don’ts, examples, and user content. By systematically defining the task scope, providing explicit operational guidelines, specifying behavioral constraints, illustrating desired results through examples, and incorporating user specific contextual information, the TIDD-EC prompt framework contributed to generating responses that are both accurate and relevant. We maintain MLLMs were instructed to select affects exclusively from the Affect Table introduced in Sect. Affect Table, rather than producing words freely. Each selected affect was required to be accompanied by a corresponding confidence score. (Appendix D) summarizes the detailed prompt configurations applied to the affect extraction task under the TIDD-EC prompt framework.
Affective Result Fusion
To mitigate potential biases inherent in individual MLLM and fully leverage the complementary strengths of different models, we integrated the results from the five models introduced in the beginning of Sect. Affect Table using the RRF algorithm (Clarke et al. 2008). RRF algorithm is an effective fusion technique that requires no training and does not depend on initial scores. This fusion strategy has been widely demonstrated to perform robustly across multimodal and retrieval tasks, effectively balancing precision and recall while enhancing the stability and consistency of final affective predictions (Samuel et al. 2025). The fusion process was explained in Figure 4.
$\text{s}(A)=\sum_{i=1}^{5}\frac{C_{i(A)}}{k+r_{i(A)}}$ (5)
User Study
In this section, we elaborated on the detailed design, implementation process, participant composition, data collection methods, and result analysis of the practical case study, aiming to verify the usability and accuracy of the InfoAffect dataset. Representative examples and the corresponding analytics for this section were illustrated in Figure 5.
Figure 5: Visual comparison of CACI results between InfoAffect dataset and crowdsourced ratings. Panels (a–i) were the representative examples cited in this section (labels on each panel); they illustrate typical cases rather than the full dataset. Each row corresponds to one affective category, indicated by colored legend markers: $\blacksquare$ Negative, $\blacksquare$ Neutral, and $\blacksquare$ Positive. The bar charts below each panel represent the proportion of affective words from each affect category.
Participants
We conducted a user study with 12 participants who had expertise in data journalism and visualization. Participants were recruited through Zhejiang University of Technology and local industry partners, which comprised two university faculty members (E1–E2) who teach courses on data visualization, three students (P1–P3), and seven industry practitioners (P4–P10), including data analysts, open-data officers, policy researchers, data consultants, marketing analyst, independent designers and design blogger. (Appendix E) summarizes the participant demographics, including gender, typical usage scenarios, and age. Participants ranged from 26 to 33 years old (mean = 29.3) and included five females and seven males. This user study received approval from the Institutional Review Board (IRB) of our university, all participants provided informed consent, and each received \$15 compensation.
Exp. I : Evaluating Usability of InfoAffect dataset
Study Procedure
Following the same processing pipeline used in constructing the InfoAffect dataset described in Sect. Affect Extraction with MLLMs, we analyzed each infographic to extract its affects. The pipeline automatically extracted affects, which were then compared with the participants’ original intent. After that, We presented the extracted affects to the participants and conducted a semi-structured interview. The specific questions and answers were provided in (Appendix F). Interviews were conducted with participants in a quiet setting, audio-recorded with informed consent, and supplemented by contemporaneous field notes capturing key quotations, clarifications, and contextual observations. Our analysis focused on how participants interpreted the detected affects, the reasoning they offered for those interpretations.
Intent vs. Observation
We contrasted each participant’s original intent for their infographic with the affect extracted by the InfoAffect dataset construction pipeline. We then discussed these contrasts in semi-structured interviews. Representative examples are shown in Figure 5. The contrasts fall into two patterns:
Validating
[…]
Correcting
The correction aspect focuses on situations where discrepancies arise between the creator’s intended affect and the extracted affects. Instead of directly modifying the design, the dataset provides an affective reference that helps creators recognize potential inconsistencies between their intended and extracted affects. Participants are then encouraged to adjust their own works—either by reinforcing insufficient affects or by reducing unintended ones.
[…]
Exp. II : Evaluating accuracy of InfoAffect dataset
This section introduced the Exp.II, which aimed to quantitatively assess the accuracy of affective extraction performed by InfoAffect dataset. The goal was to verify whether the InfoAffect results were consistent with human perceptions when evaluated by a general audience.
Study Procedure
Two sources of infographics were selected for this study: ❶ 430 items randomly sampled from the constructed InfoAffect dataset. ❷ 10 infographics collected in Exp I, in which discrepancies were observed between the participants’ intended affects and the results of InfoAffect. In total, 440 infographics were included in the study. We used Amazon Mechanical Turk (MTurk) to gather responses from a large and diverse pool of participants. 32 participants were permitted to complete this task after a 5-minute standardized introduction. We paid each participant \$1.5 for labeling 10 infographics. Additionally, we included the 12 participants from Exp I, yielding a total sample of 44 unique participants. Each infographic was presented to participants for affective evaluation. They were asked to annotate affects and confidence for each infographic.
CACI Analysis
To evaluate the agreement between InfoAffect results and human perceptions, we proposed a Composite Affect Consistency Index (CACI) that combines the dominant affect agreement and precision of top-ranked affects.For each infographic $i$, we first compute two indicators: the dominant class match $M_{i}$ and the retrieval success $S_{i}(k)$ at top-$k$. The overall CACI score is then obtained by averaging the product of these two indicators across all infographics.
Formally, CACI is defined as:
$\text{CACI}(k)=\frac{1}{N}\sum_{i=1}^{N}M_{i}\times S_{i}(k)$ (6)
Where $N=440$ is the total number of infographics mentioned above, and we evaluate CACI at $k\in\{1,3,5\}$.We define the dominant class match indicator $M_{i}$ based on three-dimensional affect vectors. For each infographic $i$, let $(n_{1}^{(i)},ne_{1}^{(i)},p_{1}^{(i)})$ represent the number of raters selecting Negative, Neutral, and Positive in the InfoAffect prediction, and $(n_{2}^{(i)},ne_{2}^{(i)},p_{2}^{(i)})$ for human annotations. Then:
$M_{i}=\begin{cases}1,&\text{if }\arg\max(n_{1}^{(i)},ne_{1}^{(i)},p_{1}^{(i)})=\arg\max(n_{2}^{(i)},ne_{2}^{(i)},p_{2}^{(i)})\\ 0,&\text{otherwise}\end{cases}$ (7)
The retrieval success indicator $S_{i}(k)\in\{0,1\}$ represents whether the top-ranked affect from human annotations appears in the top-$k$ predictions from InfoAffect. Let $a_{i}^{*}$ denote the affect word with the highest confidence from the crowdsource results for infographic $i$, and let $\text{Top}_{k}^{(i)}$ denote the set of top-$k$ predictions from InfoAffect for infographic $i$. Then $S_{i}(k)$ is computed as:
$S_{i}(k)=\begin{cases}1,&\text{if }a_{i}^{*}\in\text{Top}_{k}^{(i)}\\ 0,&\text{otherwise}\end{cases}$ (8)
CACI effectively captures both the alignment of dominant affects and the precision of top-ranked affects, providing a comprehensive measure of affective consistency between InfoAffect results and human perceptions. The multiplicative formulation in Equation 6 ensures that both aspects must be satisfied for a high score: the dominant class match indicator $M_{i}$ acts as a binary gate that enforces overall affective alignment, while the retrieval success indicator $S_{i}(k)$ measures the accuracy of affect extraction. Notably, when the predicted dominant affect class (Positive, Neutral, or Negative) does not match human perceptions, $M_{i}$ becomes 0, which drives the contribution of that sample to 0 regardless of the value of $S_{i}(k)$. Conversely, when $M_{i}=1$, the sample’s contribution is determined solely by $S_{i}(k)$.
As shown in Figure 5, we present infographic samples together with their corresponding CACI values. Table 2 presents the CACI results for with and without the Affect Table at different top-$k$. Each row shows $\text{CACI}(k)$ computed by averaging $M_{i}\times S_{i}(k)$ across all 440 infographics. The results show that the resulting $\text{CACI}_{\textit{InfoAffect}}=0.608$ (averaged across $k\in\{1,3,5\}$), and incorporating the Affect Table improved affect extraction by 13.4% on average compared to the condition without the Affect Table. These results collectively demonstrate that the affective extraction maintains high accuracy in identifying affects of infographics.
Table 2: CACI results for InfoAffect at different top-$k$. Each row shows $\mathrm{CACI}(k)$ computed by averaging $M_{i}\times S_{i}(k)$ across all 440 infographics.
Evaluation Without Affect Table With Affect Table
$\mathrm{CACI}(1)$ 0.372 0.454 (+22.0%)
$\mathrm{CACI}(3)$ 0.573 0.643 (+12.2%)
$\mathrm{CACI}(5)$ 0.664 0.726 (+9.3%)
Average 0.536 0.608 (+13.4%)
Limitations and Future Work
While InfoAffect dataset advanced affective analysis of infographics, we acknowledged two important limitations. First, our dataset size remains modest. Second, while the first-level classification results are balanced, the second-level classification results are long-tailed due to the fine-grained classification and the use of real-world infographics. In future work, we will study how specific visual and textual elements in infographics shape perceived affect by varying one factor at a time on existing designs and noting the affect shifts.
Conclusion
We presented InfoAffect dataset, a 3.5k-sample dual-modal that pairs real-world infographics with accompanied text and constrains affect labels with the Affect Table. The corpus covers six fields and is prepared with three methods to ensure alignment and quality. We extract affects with five state-of-the-art MLLMs and fuse their outputs using RRF to obtain robust affective results. We validated the resource with two experiments that assessed usability and accuracy of InfoAffect dataset. The accuracy study introduced the CACI and achieved an overall score of 0.608. InfoAffect dataset establishes a foundation for future research on affective understanding, model training, and affectively informed visualization design.
Acknowledgments
Appendix A Dataset Preparation Prompts
Task: You are given an infographic image. Your task is to produce a concise, factual, and content-faithful textual description of the infographic.
Guidelines:
Focus strictly on the information conveyed by the infographic, including the main topic, key variables, data relationships, trends, comparisons, or notable patterns.
Describe what the infographic communicates, not how it looks. Avoid stylistic, aesthetic, or subjective language (e.g., colors, layout beauty, visual appeal).
Do not speculate beyond the visible content. Do not infer causes, motivations, or implications unless they are explicitly stated in the infographic.
If numerical values, time ranges, categories, or entities are present, summarize them accurately at a high level without exhaustively listing all values.
Write in a neutral, encyclopedic tone, similar to a figure caption or news standfirst.
The description should be self-contained, understandable without seeing the image, and no longer than 3–5 sentences.
Output: A single paragraph describing the infographic’s communicative intent and key information.
Appendix B Affect Table
Results of the two-level classification. The left column corresponds to the first-level classification results, the middle column contains the second-level classification results, and the right column lists the affects initially generated by GPT-4o.
First-level Second-level Affects
Positive rise growth, grew, growing, expansion, expanding, rise, risen, rising, climb, ascending, expansive, expanded
soaring soared, soaring, surge, surged, surging, skyrocketed, booming, boost, explosive, buoyant, skyrocket, boosted
success success, successful, achievement, achieving, accomplished, triumph, victory, milestone, crowned, triumphant, victorious, dominant, dominance, confident, confidence, ambitious, ambition, competitive, competitiveness
excellent excellent, superior, exceptional, outstanding, top, highest, best, elite, premium, perfect, impressive, impressively, perfectly
innovative innovative, groundbreaking, pioneering, trailblazer, revolutionary, visionary, modern, futuristic, progressive, progressively, modernized, transformative, transformatively
effective effective, efficient, productive, practical, useful, valuable, functional, proactive, resourceful, informative, informatively, professional, professionally, strategic, strategically, practically
reliable trust, trusted, trustworthy, reliable, authentic, stable, stability, secure, security, credible, credibility, authenticity, resilient, resiliently, significant, significantly
support support, supportive, helpful, aid, assistance, collaboration, cooperative, partnership, united, encouraging, encouragingly
joy happy, happiness, joyful, jubilant, cheerful, elated, exuberant, delightful, pleasant, sweet, joy, optimistic, optimistically, hopeful, hopefully, positive, positively, enthusiastic, enthusiasm, playful, playfulness, exciting, excitingly, celebratory, celebratorily, reassuring, reassuringly
energetic vibrant, active, dynamic, energetic, lively, attractive, captivating, stunning, vivid, vibrantly, actively, dynamically, energetically, influential, influence
Neutral anticipated likely, probable, expected, projected, forecast, forecasting, potential, predicted, anticipated, anticipation, anticipatory
transitional change, changing, shift, shifted, shifting, transition, transitional, transforming, variation, modified, varied, modifying, surprising, surprisingly
pending waiting, pending, postponed, temporary, transient, interim, provisional, paused, delayed, holding
formal official, regulated, regulatory, administrative, controlled, standard, formal, institutional, legal, mandated, officially, formally, objective, objectively
practical economic, financial, commercial, industrial, corporate, market, trade, monetary, fiscal, economically, financially, commercially, expensive, expensively
statistical measured, calculated, statistical, computed, data, metric, empirical, quantified, statistically, empirically, analytical, analytically, factual, factually, comparative, comparatively, insightful, insightfully
perceived observed, perceived, perception, tracked, tracking, monitored, phenomenon, identified, detected, revealing, revealingly
mixed mixed, varied, various, miscellaneous, diverse, multiple, random, differing, diversity, disparate, disparity, divisive, divisively, controversial, controversially
normal normal, normalized, standard, habitual, routine, constant, typical, regular, common, typically, regularly, commonly, neutral, neutrally
Negative declining decline, declining, decrease, decreased, drop, dropped, falling, shrink, shrinking, downward, declined, decreasing, dropping, fallen, shrunk, disappointing, disappointingly
crashing plummet, plummeted, plunge, plunging, collapse, collapsed, crash, crashing, nosedived, tumble, plummeting, plunged, collapsing, crashed, tumbled, tumbling
fear scared, fear, fears, fearful, frightened, terrified, panic, panicked, apprehensive, dread, scary, fearfully, terrifying, panicking, apprehensively, dreadful, distressing, distressingly, somber, somberly, grave, gravely, anxious, anxiously, tragic, tragically, overwhelming, overwhelmingly
alarm alarm, alarmed, alarming, warning, emergency, threatening, threat, danger, hazardous, alarmingly, warned, threateningly, threatened, dangerous, dangerously, concerning, concerningly, serious, seriously, critical, critically, urgent, urgently, sobering, soberingly, cautious, cautiously, troubling, troublingly, worrisome, worrisomely, disturbing, disturbingly
anger anger, angry, furious, infuriating, outrage, hostile, hostility, resentment, bitter, clashes, angrily, furiously, infuriatingly, outraged, outrageous, resentful, bitterly, clashing, tense, tensely, restrictive, restrictively, frustrating, frustratingly, divided, dividedly
struggle struggle, struggles, struggling, hardship, difficulty, difficulties, tough, challenging, gruelling, struggled, difficult, challenge, problematic, problematically
pressured pressure, pressured, pressuring, strained, strains, stretched, burden, burdened, constrained, stretching, burdening, constraining
harmful damage, damaged, damaging, harm, harmful, detrimental, toxic, destructive, devastated, catastrophic, harmfully, detrimentally, toxically, destructively, devastating, catastrophically, disheartening, dishearteningly
failed fail, failed, failing, failure, defeat, defeated, unsuccessful, setback, derailed, mistake, defeating, unsuccessfully, setbacks, mistakes, mistaken
shortage shortage, scarcity, scarce, insufficient, inadequate, lack, missing, deficit, void, empty, shortages, insufficiently, inadequately, lacking, deficits, emptied
uncertain uncertain, uncertainty, unclear, unsure, skepticism, dubious, suspicion, ambiguous, tentative, uncertainties, unclearly, skeptical, skeptically, dubiously, suspicious, suspiciously, ambiguously, tentatively
lonely isolated, isolation, abandoned, abandonment, alone, lonely, excluded, alienation, vanished, exclusion, vanishing
Appendix C JSON schema
@@
The table provided a detailed explanation of the JSON schema adopted in InfoAffect dataset for organizing affective annotations. Each key–value pair specified how the detected affects, corresponding confidence scores, and relevant lexical information are stored and accessed. This standardized schema supported transparent data management, reproducibility of experiments, and seamless integration with downstream analytical workflows.
@@
Appendix D TIDD-EC Prompt Framework
Table 3 presented the detailed configuration of the TIDD-EC prompt framework employed for affect extraction. Each component defined a specific role in guiding multimodal large language models, from specifying the analytical task and procedural instructions to providing explicit behavioral constraints and contextual examples.
Table 3: Detailed configuration of the TIDD-EC prompt framework for affect extraction.
@@ Appendix D TIDD-EC Prompt Framework
D-Do Specify the actions that MLLMs should implement during the response process. Provide your analysis in the following JSON schema. Output ONLY a single valid JSON object.
@@
Appendix E Participant Information
Table 4 details the demographic and occupational composition of participants who took part in the InfoAffect dataset evaluation study. It lists their professional identities, gender, usage scenarios, and age. The group comprises researchers, students, designers, and data professionals with backgrounds in visualization and communication design.
Table 4: Participants in user study.
@@ Appendix E Participant Information
Appendix F Interview transcripts
eg.1 E1
<original intent>: I designed this infographic to highlight China’s global leadership in solar photovoltaic deployment using clear quantitative evidence. My primary intention was to convey pride and encouragement toward renewable energy advancement, emphasizing progress and global responsibility rather than competition.
<extracted affects>: Optimistic, confident, and assertive.
@@ Appendix F Interview transcripts
Q1: Do you think the affect identified aligns with the affect youoriginally intended to convey in this infographic?
A1: Largely, yes. The system’s identification of “optimistic” and “confident” aligns well with my goal of promoting a positive outlook on renewable energy leadership.
@@
Q2: If the extracted affects differs from your original intent, how would you describe that difference?
A2: [The extracted affect aligns with the original intent.]
@@
Q3: Do you think the visual elements, such as color, composition, or typography, contributed to this unintended affects?
A3: [The extracted affect aligns with the original intent.]
@@
Q4: Does the analysis results help you better understand how affect is represented in your visual design?
A4: Definitely. It helped me realize how small visual decisions, like using red in a large proportion or positioning the label prominently, can shift the perceived affect from informative to persuasive.